If the demand curve of a profit maximizing monopolist is given as Q = 40 − 0.2P and cost function as C = 30 + 30Q, find equilibrium output level, monopolist price, and profit


1
Expert's answer
2021-06-30T11:50:16-0400

Given

"Q = 40 − 0.2P\\C = 30 + 30Q"

For profit monopolist ,the equilibrium output is determined at point where marginal revenue is equal to marginal cost.

"TR=P×Q\\Q=40-0.2P\\0.2p =40-Q\\P=200-5Q"


"TR=(200-5Q)Q\\TR=200Q-5Q^2\\MR=200-10Q......(1)"


"TC=30+30Q\\MC=30......(2)"

At equilibrium

"MR=MC\\200-10Q=30\\170=10Q\\Q=17"

Substitute Q to get value of P

"P=200-5Q\\=200-5(17)\\P=115"


Profit=Total revenue-Total cost

"=(P×Q)-(TC)\\=(200-5Q)Q-(30+30Q)\\=200Q-5Q^2-30+30Q\\=230Q-5Q^2-30"

Substitute for Q=17

Profit

"=230×17-5(17)^2-30\\=3910-1445-30\\=2435"